# Nearly-Optimal Bandit Learning in Stackelberg Games with Side Information

### v1

@@ 1 Introduction
Harris et al. [14] formalize this setting and provide online learning algorithms for the leader when the followers and contextual information change over time. Their algorithms obtain $\tilde{O}(T^{1/2})$ regret [^2] Regret is the cumulative difference between the highest possible cumulative utility and the algorithm’s cumulative utility. under full feedback (i.e. when information about the follower is revealed to the leader after each round) where $T$ is the number of time-steps, but only $\tilde{O}(T^{2/3})$ regret under the more challenging (and more realistic) bandit feedback setting, where only the follower’s action is revealed.
@@
We close the gap from $\tilde{O}(T^{2/3})$ to $\tilde{O}(T^{1/2})$ regret under bandit feedback, which matches known lower bounds up to logarithmic factors. As in Harris et al. [14], we study two settings: one in which the sequence of contextual information is chosen adversarially and the sequence of followers is chosen stochastically, and the setting where the contextual information is chosen stochastically and the followers are chosen adversarially. Moreover, the algorithms of Harris et al. [14] are not applicable when the follower’s utility depends on the contextual information, an assumption which we do not need.
@@
Algorithm 1 may be thought of as transforming the leader’s learning problem into a different space in which the rewards are linear, then playing a contextual bandit algorithm over a finite set of actions in this alternative space. While the set of actions in both spaces is exponentially large, we obtain regret guarantees which only depend polynomially on the size of the problem by working in the dual space. In Section 4 we show how to apply our algorithm to learning in other settings which exhibit the same type of structure; specifically (i) learning in second-price auctions with side information and (ii) online Bayesian persuasion with side information. We are the first to study either of these settings, to the best of our knowledge, despite the fact that side information naturally arises in both auctions and Bayesian persuasion settings. Our results largely carry over to these applications as-is, although we need to discretize the learner’s action space in a different way than we do for Stackelberg games. Finally, we empirically evaluate Algorithm 1 in Section 5 and highlight directions for future research in Section 6.
@@ 1.1 Related Work
Conitzer and Sandholm [7] provide algorithms and prove NP-Hardness results for the problem of computing equilibrium in various Stackelberg game settings when all parameters of the problem are known. A line of work on learning Stackelberg games [17, 20, 2] relaxes the assumption that all parameters of the problem are known to the leader, and instead posits that they are given a number of (player actions, outcome) tuples to learn from.
Our work falls under the category of online learning in Stackelberg games, where the sequence of data arrives sequentially instead of all at once. This setting was first introduced by Balcan et al. [3] and was generalized to handle settings with side information in Harris et al. [14].
Other recent work on learning in Stackelberg games includes learning in cooperative Stackelberg games (e.g. Zhao et al. [26], Donahue et al. [10]), strategizing against a follower who plays a no-regret learning algorithm (e.g. Braverman et al. [5], Deng et al. [9]), and learning various structured Stackelberg games such as strategic classification (e.g. Hardt et al. [13], Dong et al. [11]) and performative prediction (e.g. Perdomo et al. [21], Hardt and Mendler-Dünner [12]).
@@
Our algorithms are also applicable to generalizations of the problems of online learning in simultaneous second-price auctions [8] and online Bayesian persuasion [6].
Bernasconi et al. [4] use a similar reduction to obtain $\tilde{O}(T^{1/2})$ regret in online Bayesian persuasion, learning in auctions, and learning in Stackelberg games without side information. While their main result is a reduction to adversarial linear bandits with a constant (and possibly infinite) action set, ours is a reduction to linear contextual bandits with time-varying (but finite) action sets. As such, more care needs to be taken to show that a contextual bandit algorithm satisfies the requirements to be used by our reduction. Additionally, while both reductions leverage the linear structure which is induced from having finitely-many follower types, our extension to unknown leader utilities in Section 3.3 uses a more general version of this linear structure in order to compensate for the additional uncertainty from unknown utilities.
@@
Finally, one may view our setting as a special type of contextual bandit problem with continuous action spaces and non-linear rewards. While one could, in principle, attempt to apply a black-box contextual bandit algorithm to our setting (e.g. Syrgkanis et al. [24, 25], Rakhlin and Sridharan [22]), we are not aware of any algorithms which obtain meaningful performance guarantees under this reward structure without (1) making additional assumptions about the learner’s knowledge of the sequence of contexts they will face and (2) obtaining generally worse rates.
@@ 2 Preliminaries
We use $\Delta(\mathcal{A})$ to denote the probability simplex over the (finite) set $\mathcal{A}$, and $[N]:=\{1,\ldots,N\}$ to denote the set of integers from $1$ to $N\in\mathbb{N}_{>0}$.
@@
We assume that the follower in each round is one of $K<\infty$ types $f_{t}\in\{\alpha^{(1)},\ldots,\alpha^{(K)}\}$, where follower type $\alpha^{(i)}$ corresponds to utility function $u_{\alpha^{(i)}}$. We assume that $u_{\alpha^{(1)}},\ldots,u_{\alpha^{(K)}}$ are known to the leader, but the identity of follower $f_{t}$ is never revealed. [^3] Balcan et al. [3] show that learning is impossible when $K=\infty$ in (non-contextual) Stackelberg games, which implies an impossibility result for our setting. This setting is referred to as bandit feedback in the literature on online learning in Stackelberg games [3, 14].
Given a context $\mathbf{z}_{t}$ and leader mixed strategy $\mathbf{x}_{t}$, follower $f_{t}$’s best-response is
$b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}):=\arg\max_{a_{f}\in\mathcal{A}_{f}}u_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t},a_{f}),$
where ties are broken in a fixed-but-arbitrary way.
We measure the performance of the leader via the notion of regret:
@@
The leader’s contextual Stackelberg regret with respect to context sequence $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ and follower sequence $f_{1},\ldots,f_{T}$ is
$R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\mathbf{x}_{t}^{*},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}^{*}))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$
where
$\mathbf{x}_{t}^{*}=\pi^{*}(\mathbf{z}_{t}):=\arg\max_{\mathbf{x}\in\Delta(\mathcal{A}_{l})}\sum_{\tau:\mathbf{z}_{\tau}=\mathbf{z}_{t}}u(\mathbf{z}_{\tau},\mathbf{x},b_{f_{\tau}}(\mathbf{z}_{\tau},\mathbf{x}))$
is the mixed strategy played by the optimal-in-hindsight policy $\pi^{*}$ at time $t$.
Since previous work [14] shows that no-regret learning is impossible (i.e. there exists no algorithm for which $R(T)=o(T)$) when the sequence of contexts and the sequence of followers are chosen jointly by an adversary with knowledge of the leader’s algorithm, we focus on two natural relaxations: the setting where the sequence of contexts is chosen by an adversary and the sequence of follower types are drawn from an unknown (stationary) distribution (Section 3.1) and the setting where the sequence of follower types are chosen by an adversary and the sequence of contexts are drawn from an unknown distribution (Section 3.2). All of our results are applicable to the simpler setting where both the contexts and follower types are chosen stochastically.
@@ 3 A Reduction to Linear Contextual Bandits
      Play action $a_{l,t}\sim\mathbf{x}_{t}$
      Call .$\mathcal{R}.\mathrm{observeUtility}(\mathbf{v}_{t},u(\mathbf{z}_{t},a_{l,t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})))$
@@
Our main result is an algorithm (Algorithm 1) that achieves $\tilde{O}(T^{1/2})$ regret in both the setting where contexts are chosen adversarially and follower types are chosen stochastically (Section 3.1) and the setting where the contexts are chosen stochastically and follower types are chosen adversarially (Section 3.2). While the leader’s utility is a non-linear function of their mixed strategy $\mathbf{x}_{t}$ in any given round (due to the follower’s best-response $b(\mathbf{z}_{t},\mathbf{x}_{t})$), we can “linearize” the problem by leveraging the fact that the leader’s utility can be written as $u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))=\langle\mathbf{u}(\mathbf{z}_{t},\mathbf{x}_{t}),\mathbf{1}_{f_{t}}\rangle$, where
$\mathbf{u}(\mathbf{z},\mathbf{x}):=\begin{bmatrix}u(\mathbf{z},\mathbf{x},b_{\alpha^{(1)}}(\mathbf{z},\mathbf{x}))\\ \vdots\\ u(\mathbf{z},\mathbf{x},b_{\alpha^{(K)}}(\mathbf{z},\mathbf{x}))\end{bmatrix}\in\mathbb{R}^{K}$
is the vector of utilities the leader would receive against each follower type given context $\mathbf{z}$ and mixed strategy $\mathbf{x}$, and $\mathbf{1}_{f_{t}}\in\mathbb{R}^{K}$ is a one-hot vector with a $1$ in the $f_{t}$-th component and zeros elsewhere. Since $u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f}(\mathbf{z}_{t},\mathbf{x}_{t}))$ is a linear function of $\mathbf{u}(\mathbf{z}_{t},\mathbf{x}_{t})$, one can use an off-the-shelf linear contextual bandit algorithm to pick a vector $\mathbf{v}_{t}$ in the image of $\mathbf{u}(\mathbf{z}_{t},\cdot)$, then invert the mapping to find the mixed strategy $\mathbf{x}_{t}$ such that $\mathbf{v}_{t}=\mathbf{u}(\mathbf{z}_{t},\mathbf{x}_{t})$.
Algorithm 1 takes as input a linear contextual bandit algorithm $\mathcal{R}$, which, (1) when given a (finite) set of actions $U_{t}$, returns an element $\mathbf{v}_{t}\in U_{t}$ (.$\mathcal{R}.\mathrm{recommend}()$) and (2) updates its internal parameters when given an action $\mathbf{v}_{t}$ and a realized utility $u_{t}\in[-1,1]$ (.$\mathcal{R}.\mathrm{observeUtility}()$). Finally, while the leader’s action space $\Delta(\mathcal{A}_{l})$ is infinitely large (and thus, so is the dual space $\tilde{U}_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{x})\;:\;\mathbf{x}\in\Delta(\mathcal{A}_{l})\}$), the leader incurs essentially no loss in utility by restricting themselves to a finite (but exponentially-large) set of context-dependent points $\mathcal{E}_{t}$ (defined in Section 3.4), which roughly correspond to the set of extreme points of convex polytopes which are induced by the followers’ best-responses. [^4] More details on how $\mathcal{E}_{t}$ is defined may be found in Section 3.4. As such, our algorithm operates on the set of utility vectors $U_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{x})\;:\;\mathbf{x}\in\mathcal{E}_{t}\}$ in each round. [^5] This is important, as the regret minimizers we instantiate Algorithm 1 with in Section 3.1 and Section 3.2 both require the action set to be finite.

### v2

@@ 1 Introduction
Harris et al. (2024) formalize this setting and provide online learning algorithms for the leader when the followers and contextual information change over time. Their algorithms obtain $\tilde{O}(T^{1/2})$ regret [^1] Regret is the cumulative difference between the best possible policy’s utility and the leader’s utility. under full feedback (i.e. when information about the follower is revealed to the leader after each round) where $T$ is the number of time-steps, but only $\tilde{O}(T^{2/3})$ regret under the more challenging (and more realistic) bandit feedback setting, where only the follower’s action is revealed.
@@
We close the gap from $\tilde{O}(T^{2/3})$ to $\tilde{O}(T^{1/2})$ regret under bandit feedback, which matches known lower bounds up to logarithmic factors. As in Harris et al. (2024), we study two settings: one in which the sequence of contextual information is chosen adversarially and the sequence of followers is chosen stochastically, and the setting where the contextual information is chosen stochastically and the followers are chosen adversarially. Moreover, the algorithms of Harris et al. (2024) are not applicable when the follower’s utility depends on the contextual information, an assumption which we do not need.
@@
In Section 4 we show how to apply our algorithm to learning in other settings which exhibit the same type of structure; specifically (i) learning in second-price auctions with side information and (ii) online Bayesian persuasion with side information. We are the first to study either of these settings, to the best of our knowledge, despite the fact that side information naturally arises in both auctions and Bayesian persuasion settings. Our results largely carry over to these applications as-is, although we need to discretize the learner’s action space in a different way than we do for Stackelberg games.
Our work is conceptually related to Bernasconi et al. (2023), who use a similar reduction to obtain $\tilde{O}(T^{1/2})$ regret in online Bayesian persuasion, learning in auctions, and learning in Stackelberg games without side information. While their main result is a reduction to adversarial linear bandits, our problem reduces to a linear contextual bandit problem with an infinite action set. Furthermore, either the action set or the sequence of contexts will be chosen adversarially, depending on which setting we are in. While such contextual bandit problems are generally intractable, we leverage the special structure present in our setting to obtain positive results. In particular, we observe that the optimal strategy at each time-step will always belong to a time-varying, but finite, set. Therefore, we can discretize the utility space in a way that allows us to apply our reduction with two specific linear contextual bandit algorithms that allow for both time-varying action sets and adversarially-chosen contexts/action sets. Additionally, while both our reduction and theirs leverage the linear structure which is induced from having finitely-many follower types, our extension to unknown leader utilities in Section 3.3 uses a more general version of this linear structure in order to compensate for the additional uncertainty from unknown utilities. Our reduction is also more streamlined, as it does not require a per-round application of Caratheodory’s Theorem, due to our discretization step.
@@ 1.1 Related work
Conitzer and Sandholm (2006) provide algorithms and prove NP-Hardness results for the problem of computing equilibrium in various Stackelberg game settings when all parameters of the problem are known. A line of work on learning Stackelberg games Letchford et al. (2009); Peng et al. (2019); Bacchiocchi et al. (2025) relaxes the assumption that all parameters of the problem are known to the leader, and instead posits that they are given a number of (player actions, outcome) tuples to learn from.
Our work falls under the category of online learning in Stackelberg games, where the sequence of data arrives sequentially instead of all at once. This setting was first introduced by Balcan et al. (2015) and was generalized to handle settings with side information in Harris et al. (2024). Other recent work on learning in Stackelberg games includes learning in cooperative Stackelberg games (e.g. Zhao et al. (2023); Donahue et al. (2024)), strategizing against a follower who plays a no-regret learning algorithm (e.g. Braverman et al. (2018); Deng et al. (2019)), and learning various structured Stackelberg games such as strategic classification (e.g. Hardt et al. (2016); Dong et al. (2018)), performative prediction (e.g. Perdomo et al. (2020); Hardt and Mendler-Dünner (2023)), and principal-agent problems (e.g. Ho et al. (2016)).
Online learning in Stackelberg games is conceptually related to the online optimization of piecewise Lipschitz functions, as the underlying reward function after fixing any fixed follower type (and piece of side information) is piecewise linear Dick et al. (2020); Sharma et al. (2020); Balcan et al. (2018). However the techniques used in this line of work are not applicable to our setting, since the follower type and contextual information change from round-to-round.
@@
Our algorithms are also applicable to generalizations of the problems of online learning in simultaneous second-price auctions Daskalakis and Syrgkanis (2022) and online Bayesian persuasion Castiglioni et al. (2020). Flajolet and Jaillet (2017) also study online learning in second-price auctions with side information. They consider single-item auction settings with a budget constraint, while we consider combinatorial auctions with no budget constraints. We also consider a more general form of adversarial feedback than they consider.
@@
Finally, one may view our setting as a special type of contextual bandit problem with continuous action spaces and non-linear rewards. While one could, in principle, attempt to apply a black-box contextual bandit algorithm to our setting (e.g. Syrgkanis et al. (2016a; b); Rakhlin and Sridharan (2016)), we are not aware of any algorithms which obtain meaningful performance guarantees under this reward structure without (1) making additional assumptions about the learner’s knowledge of the sequence of contexts they will face and (2) obtaining generally worse rates.
@@ 2 Preliminaries
We use $\Delta(\mathcal{A})$ to denote the probability simplex over the (finite) set $\mathcal{A}$, and $N]:=\{1,\ldots,N\}$ to denote the set of integers from $1$ to $N\in\mathbb{N}_{>0}$.
@@
We assume that the follower in each round is one of $K<\infty$ types $f_{t}\in[K]$, where follower type $i\in[K]$ corresponds to utility function $u_{i}$. We assume that $u_{1},\ldots,u_{K}$ are known to the leader, but the identity of follower $f_{t}$ is never revealed. [^2] Balcan et al. (2015) show that learning is impossible when $K=\infty$ in (non-contextual) Stackelberg games, which implies an impossibility result for our setting. This setting is referred to as bandit feedback in the literature on online learning in Stackelberg games Balcan et al. (2015); Harris et al. (2024).
Given a context $\mathbf{z}_{t}$ and leader mixed strategy $\mathbf{x}_{t}$, follower $f_{t}$’s best-response is $b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}):=\arg\max_{a_{f}\in\mathcal{A}_{f}}u_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t},a_{f})$, where ties are broken in an unknown-but-fixed way. We measure the performance of the leader via the notion of regret:
@@
The leader’s contextual Stackelberg regret with respect to context sequence $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ and follower sequence $f_{1},\ldots,f_{T}$ is $R(T):=\sum_{t=1}^{T}u(\mathbf{z}_{t},\mathbf{x}_{t}^{*},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}^{*}))-u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))$, where $\mathbf{x}_{t}^{*}=\pi^{*}(\mathbf{z}_{t}):=\arg\max_{\mathbf{x}\in\Delta(\mathcal{A}_{l})}\sum_{\tau:\mathbf{z}_{\tau}=\mathbf{z}_{t}}u(\mathbf{z}_{\tau},\mathbf{x},b_{f_{\tau}}(\mathbf{z}_{\tau},\mathbf{x}))$ is the mixed strategy played by the optimal-in-hindsight policy $\pi^{*}$ at time $t$.
Since previous work Harris et al. (2024) shows that no-regret learning is impossible (i.e. there exists no algorithm for which $R(T)=o(T)$) when the sequence of contexts and the sequence of followers are chosen jointly by an adversary with knowledge of the leader’s algorithm, we focus on two natural relaxations: the setting where the sequence of contexts is chosen by an adversary and the sequence of follower types are drawn from an unknown (stationary) distribution (Section 3.1) and the setting where the sequence of follower types are chosen by an adversary and the sequence of contexts are drawn from an unknown distribution (Section 3.2). All of our results are applicable to the simpler setting where both the contexts and follower types are chosen stochastically.
While the leader’s action space $\Delta(\mathcal{A}_{l})$ is infinitely large, we follow the lead of previous work and consider a discretization which is nearly wothout loss of generality. The following two definitions are from Harris et al. (2024).
Definition 2.2 (Contextual Follower Best-Response Region).
For follower type $i\in[K]$, follower action $a_{f}\in\mathcal{A}_{f}$, and context $\mathbf{z}\in\mathcal{Z}$, let $\mathcal{X}_{\mathbf{z}}(i,a_{f})\subseteq\Delta(\mathcal{A}_{l})$ denote the set of all leader mixed strategies such that a follower of type $i$ best-responds to all $\mathbf{x}\in\mathcal{X}_{\mathbf{z}}(i,a_{f})$ by playing action $a_{f}$ under context $\mathbf{z}$, i.e., $\mathcal{X}_{\mathbf{z}}(i,a_{f})=\{\mathbf{x}\in\mathcal{X}:b_{i}(\mathbf{z},\mathbf{x})=a_{f}\}.$
Definition 2.3 (Contextual Best-Response Region).
For a given function $\sigma:[K]\rightarrow\mathcal{A}_{f}$, let $\mathcal{X}_{\mathbf{z}}(\sigma)$ denote the set of all leader mixed strategies such that under context $\mathbf{z}$, a follower of type $i$ plays action $\sigma(i)$ for all $i\in[K]$, i.e. $\mathcal{X}_{\mathbf{z}}(\sigma)=\cap_{i\in[K]}\mathcal{X}_{\mathbf{z}}(i,\sigma(i)).$
Definition 2.4 ($\delta$-approximate extreme points).
[…]
Balcan et al. (2015) show that $|\mathcal{E}_{t}|=O((KA_{f}^{2})^{A_{l}}A_{f}^{K})$. By Lemma 4.4 in Harris et al. (2024), restricting the learner to policies which only play strategies in $\mathcal{E}_{t}$ at round $t$ leads to at most $O(1)$ additional regret. We will use fact throughout the sequel.
@@ 3 A reduction to linear contextual bandits
   Play action $a_{l,t}\sim\mathbf{x}_{t}$ and call $\mathcal{R}.\mathrm{observeUtility}(\mathbf{v}_{t},u(\mathbf{z}_{t},a_{l,t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t})))$
@@
Our main result is an algorithm (Algorithm 1) that achieves $\tilde{O}(T^{1/2})$ regret in both the setting where contexts are chosen adversarially and follower types are chosen stochastically (Section 3.1) and the setting where the contexts are chosen stochastically and follower types are chosen adversarially (Section 3.2). While the leader’s utility is a non-linear function of their mixed strategy $\mathbf{x}_{t}$ in any given round (due to the follower’s best-response $b(\mathbf{z}_{t},\mathbf{x}_{t})$), we can “linearize” the problem by leveraging the fact that the leader’s utility can be written as $u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x}_{t}))=\langle\mathbf{u}(\mathbf{z}_{t},\mathbf{x}_{t}),\mathbf{1}_{f_{t}}\rangle$, where $\mathbf{u}(\mathbf{z},\mathbf{x}):=[u(\mathbf{z},\mathbf{x},b_{1}(\mathbf{z},\mathbf{x})),\dots,u(\mathbf{z},\mathbf{x},b_{K}(\mathbf{z},\mathbf{x}))]^{\top}\in\mathbb{R}^{K}$ is the vector of utilities the leader would receive against each follower type given context $\mathbf{z}$ and mixed strategy $\mathbf{x}$, and $\mathbf{1}_{f_{t}}\in\mathbb{R}^{K}$ is a one-hot vector with a $1$ in the $f_{t}$-th component and zeros elsewhere. Since $u(\mathbf{z}_{t},\mathbf{x}_{t},b_{f}(\mathbf{z}_{t},\mathbf{x}_{t}))$ is a linear function of $\mathbf{u}(\mathbf{z}_{t},\mathbf{x}_{t})$, one can use an off-the-shelf linear contextual bandit algorithm to pick a vector $\mathbf{v}_{t}$ in the image of $\mathbf{u}(\mathbf{z}_{t},\cdot)$, then invert the mapping to find the mixed strategy $\mathbf{x}_{t}$ such that $\mathbf{v}_{t}=\mathbf{u}(\mathbf{z}_{t},\mathbf{x}_{t})$.
Algorithm 1 takes as input a linear contextual bandit algorithm $\mathcal{R}$, which, (1) when given a (finite) set of actions $U_{t}$, returns an element $\mathbf{v}_{t}\in U_{t}$ ($\mathcal{R}.\mathrm{recommend}()$) and (2) updates its internal parameters when given an action $\mathbf{v}_{t}$ and a realized utility $u_{t}\in[-1,1]$ ($\mathcal{R}.\mathrm{observeUtility}()$). Finally, while the leader’s action space $\Delta(\mathcal{A}_{l})$ is infinitely large (and thus, so is the dual space $\tilde{U}_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{x})\;:\;\mathbf{x}\in\Delta(\mathcal{A}_{l})\}$), the leader incurs essentially no loss in utility by restricting themselves to a finite (but exponentially-large) set of context-dependent points $\mathcal{E}_{t}$, which roughly correspond to the set of extreme points of convex polytopes which are induced by the followers’ best-responses.As such, our algorithm operates on the set of utility vectors $U_{t}:=\{\mathbf{u}(\mathbf{z}_{t},\mathbf{x})\;:\;\mathbf{x}\in\mathcal{E}_{t}\}$ in each round. [^3] This is important, as the regret minimizers we instantiate Algorithm 1 with in Section 3.1 and Section 3.2 both require the action set to be finite.